Plates of area A are arranged as shown. The distance between them is d. Find the net capacitance
1
Expert's answer
2020-05-12T09:58:26-0400

We know that


"electric\ field ,\ E=\frac{\sigma}{ \epsilon_0}=\frac{q}{A\epsilon_0}\\voltage=V=Ed\\C=\frac{q}{V}"

putting all values

we get

"C=\frac{AE\epsilon_0}{Ed}=\frac{\epsilon_0A}{d}"